# Large Model Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

### v1

@@ III-D Knowledge Cooperation for LM Agents
Knowledge Fusion Output fusion Directly aggregate the outputs of different models [126]
Neural network fusion Merge multiple neural networks at the parameter level [127]
Knowledge Udating External Knowledge Bases
Adjust LM via external knowledge, e.g., Internet, knowledge bases, and
human feedback
@@
Parametric Knowledge Selective fine-tuning of model’s internal parameter knowledge [129, 130, 131]
TABLE VIII: A Summary of Knowledge Searching Approaches for LM Agents
Mode Type Technology Ref.
Hallucination Reduction Use KGs to enhance factual awareness in LMs [47]
KGs Entity Disambiguation Use LMs to identify and align entities across heterogeneous KGs [132]
Data Race Context-aware concurrent fuzz test combining KGs & LMs [133]
Static Knowledge Base Retrieve knowledge from static sources for accurate reasoning [15]
RAG Dynamic Knowledge Base
Use real-time updated data sources to help LMs handle complex
and unforeseen situations
[76, 84]
Self-Guided Retrieval Use SKR to help LLM adaptively invoke external resources [71]
III-D1 Knowledge Synchronization
As summarized in Table VII, knowledge synchronization refers to the sharing and updating of knowledge among multiple LM agents to ensure consistency in decision-making, which includes knowledge transfer, knowledge fusion, and knowledge updating.
  • Knowledge transfer: It refers to transferring the implicit parameter knowledge learned by one LM agent to other LM agents. Kang et al. [123] propose an online distillation method that enhances LLMs by retrieving relevant knowledge from external knowledge bases, to generate high-quality reasoning processes. The knowledge (e.g., reasoning results) of small language models can be leveraged to improve the performance of LLMs in knowledge-intensive tasks. However, as the scale of LLMs grows, LLM training via such knowledge transfer becomes complex and computationally intensive. To address this issue, Wu et al. [124] develop a set of 2.58M instructions for fine-tuning models, generating a group of different models that maintain baseline model performance while being much smaller in size. Besides, Zhong et al. [125] propose a parametric knowledge transfer method that extracts and aligns knowledge parameters using sensitivity techniques and uses the LoRA module as an intermediary mechanism to inject this knowledge into smaller models, transferring the implicit knowledge of smaller models.
  • Knowledge fusion: It refers to integrating and optimizing knowledge from different LM agents to form a robust and comprehensive common knowledge base. Jiang et al. [126] propose an ensemble framework named LLM-blender that directly aggregates the outputs of different models to improve prediction performance and robustness. However, it requires maintaining multiple trained LLMs and executing each LLM during inference, which can be impractical for LM agents. To address this issue, Wan et al. [74] propose an implicit knowledge fusion framework that evaluates the predictive probability distributions of multiple LLMs and uses the distribution for continuous training the target model. Moreover, Wortsman et al. [127] combine multiple neural networks into a single network named “Modelsoups” through arithmetic operations at the parameter level. This method typically assumes a unified network architecture and attempts to map the weights between different neural networks.
  • Knowledge update: The timely synchronization of latest knowledge into AI models can be divided into two lines: updates to external knowledge bases and updates to internal parameter knowledge. (i) For external knowledge bases, there are three main approaches: feedback enhancement, network enhancement, and retrieval enhancement. Tandon et al. [75] pair LMs with a growing memory to train a correction model, where users identify output errors and provide general feedbacks on how to correct them. Network enhancement uses the Internet to update knowledge. Lazaridou et al. [128] use few-shot prompting to learn to adjust LMs based on the information returned from Google searches. For retrieval enhancement, Trivedi et al. [58] propose a new multi-step QA method named IRCoT to interleave retrieval with steps in CoT, which first uses CoT to guide retrieval and then uses retrieval results to improve CoT. (ii) For internal parameter knowledge, it mainly includes three methods: knowledge editing, Parameter-Efficient Fine-Tuning (PEFT), and continual learning. Knowledge editing is primarily used for quickly correcting specific errors in the model or updating a small portion of the model’s knowledge, making it suitable for fine-grained corrections. Chen et al. [129] propose a dual-stage learning algorithm called RECKONING, which folds the provided contextual knowledge into the model’s parameters and uses backpropagation to update the parameters, thereby improving the accuracy of LMs reasoning. PEFT reduces the number of parameters to be adjusted through optimization techniques for reduced computational overheads. Hu et al. [130] introduce LLM-Adapters, an easy-to-use framework that integrates various adapters into LLMs and perform these adapter-based LLM PEFT methods for different tasks. Continual learning ensures that AI model can continuously learn while receiving new tasks and new data without forgetting previously learned knowledge. Qin et al. [131] propose ELLE, which flexibly extends the breadth and depth of existing PLMs, allowing the model to continuously grow as new data flows in.
III-D2 Knowledge Searching
@@
LMs not only rely on the learned knowledge during pre-training but also can dynamically access and query external knowledge bases (e.g., databases, the Internet, and KGs) to obtain the latest information to help reasoning, as summarized in Table VIII.
① KGs: In KG, knowledge entities and their relationships are represented as structured graphs containing nodes and edges, which is widely adopted for requiring deep semantic understanding and complex relational reasoning. For QA tasks, Guan et al. [47] study a KG-enabled factual awareness enhancement method of LLMs for improved accuracy of AI models by integrating structured knowledge. KGs are also advantageous for integrating diverse information from different data sources. Zhang et al. [132] propose a fully automatic KG alignment method, using LLMs to identify and align entities in different KGs, aiming to solve the heterogeneity problem between different KGs and integrate multi-source data. Zhang et al. [133] propose a context-aware concurrent fuzz testing method that combines KGs with LLMs to effectively identify and handle data race problems in concurrent programs.
@@
② RAG: RAG technology combines information retrieval and generation models by first retrieving relevant contents and then generating answers based on these contents, suitable for fields requiring the latest information and extensive knowledge coverage. According to the data type, it can be divided into two categories. (i) RAG based on static knowledge bases, such as Wikipedia and documents. Lewis et al. [15] demonstrate how RAG generates accurate and contextually relevant responses by retrieving relevant knowledge chunks from static sources such as Wikipedia. (ii) RAG based on dynamic knowledge bases, such as news APIs, which contains two lines: exploring new knowledge or retrieving past knowledge. For new knowledge exploration, Dai et al. [76] explore the use of RAG technology for improved safety and reliability of autonomous driving systems by utilizing real-time updated data sources, including in-vehicle sensor data, traffic information, and other driving-related dynamic data. It demonstrates RAG’s potential in handling complex environments and responding to unexpected situations. For past knowledge retrieval, Kuroki et al. [84] develop a novel vector database named coordination skill database to efficiently retrieve past memories in multi-agent scenarios to adapt to new cooperation missions. By harnessing both internal and external knowledge, Wang et al. [71] propose a method called Self-Knowledge-Guided Retrieval (SKGR), which enhances retrieval by allowing LLMs to adaptively call external resources when handling new problems.
TABLE IX: A Summary of Knowledge and LM Co-driven Approaches Among LM Agents
Mode Type Technology Ref.
Pre-training
Use a explainable neuro-symbolic knowledge base, composed of vector
representations of entities and relations from existing knowledge bases
[134]
Knowledge Base-
Enhanced LMs
Fine-tuning Generate large-scale diverse prompts via numerous traversal paths in KG [73]
Inference
Combine LLMs with KGs to capture factual knowledge and update
knowledge base
[47]
Entity and
relationship extraction
Automatically extract entities and relationships [135, 79, 80]
LM-Enhanced
KGs
KG Construction
Grapher: An end-to-end multi-stage approach that generates graph nodes and
links using LMs
[136]
KG Verification Using a small LM as a verification module to validate and correct the output [83]
KG Completion Encode text or generate new facts to better use textual & semantic info [77, 137]
III-D3 Knowledge and LM Co-driven
Knowledge and LM co-driven paradigms contains two lines of approaches: knowledge base-enhanced LMs and LM-assisted KGs, as summarized in Table IX.
① Knowledge base-enhanced LMs: Knowledge base can enhance LM inference at various stages. (i) In the pre-training stages, Sun et al. [134] propose an explainable neuro-symbolic knowledge base, where the fact memory is composed of a triple of vector representations of entities and relations from existing knowledge bases. These vector representations are integrated into the LLM during pre-training. (ii) In the fine-tuning stage, considering that existing methods often overwrite the original parameters of pre-trained models when injecting knowledge, Wang et al. [73] propose the K-Adapter framework. This framework uses RoBERTa as the backbone model and assigns a neural adapter to each type of injected knowledge. These adapters can be trained effectively in a distributed manner, thereby improving the performance of LMs on specific tasks. (iii) In the inference stage, most existing methods focus on factual information related to entities explicitly mentioned in the query. Guan et al. [47] are the first to consider verification during the inference process of LLMs. They proposed a new KG-based retrofitting framework, which automatically adjusts the initial responses generated by LLMs based on factual knowledge stored in KGs, effectively improving inference accuracy.
② LM-enhanced KGs: The information extraction and knowledge fusion capabilities of LMs assist the integration and updating of diverse data during the construction and completion of KGs. (i) Entity and relationship extraction. Traditional methods typically rely on specific modules to handle each data type separately, such as text data, image data, or structured data. However, LMs can automatically extract entities and relationships. De Cao et al. [135] propose GENRE, a system that retrieves entities in knowledge graphs by generating their unique names in an autoregressive manner from left to right, which better captures the relationship between context and entity names while significantly reducing memory consumption through an encoder-decoder architecture. However, entity linking speed affects the inference speed of LMs. Ayoola et al. [79] propose ReFinED, an efficient end-to-end entity linking model that uses fine-grained entity types and descriptions for linking, achieving speeds over 60 times faster than existing methods. LMs can also extract relationships between entities. Ma et al. [80] propose DREEAM, a self-training strategy for document-level relationship extraction, which is the first system to adopt a self-training strategy for evidence retrieval, learning entity relationships from automatically generated evidence from large datasets, addressing the issues of high memory consumption and limited annotation availability. (ii) KG construction, which can be done through constructing from raw text and extracting from LLMs. Melnyk et al. [136] propose a novel end-to-end multi-stage system for efficient KG construction from text descriptions, which first utilizes LLMs to generate KG entities, followed by a simple relation construction head. (iii) KG verification. Apart from KG construction, LMs can be employed to verify KGs. Han et al. [83] propose a prompt framework for iterative verification, using small LMs to correct errors in KG generated by LLMs such as ChatGPT. (iv) KG completion, i.e., inferring missing facts in a given KG. Traditional KG completion methods merely focus on the structure of KGs, without considering extensive textual information. LLMs holds the potential to enhance KG completion performance via encoding text or generating facts. Shen et al. [77] use LLMs as encoders, primarily capturing the semantic information of knowledge graph triples through the model’s forward pass, and then reconstructing the knowledge graph’s structure by calculating a loss function, thereby better integrating semantic and structural information. As structure-specific KG construction models are mutually incompatible and cannot adapt to emerging knowledge graphs, Chen et al. [137] use LLMs as generators to propose KG-S2S, a Seq2Seq generative framework that represents knowledge graph facts as“flat” text to address the issue of heterogeneous graph structures and generate the information lost during flattening in the completion process.
IV Security Threats & Countermeasures to Large Model Agents
In this section, we present a comprehensive review of security threats related to LM agents and examine the state-of-the-art countermeasures to defend against them, by investigating the recent research advancements. Fig. 15 illustrates the taxonomy of security threats to LM agents. Firstly, the definition, categorization, causes, and corresponding countermeasures for hallucination threats are provided in Sect. IV-A. Following that, we discuss adversarial attacks including adversarial input attacks and prompt hacking attacks and review corresponding countermeasures in Sect. IV-B. Next, we present poisoning and backdoor attacks to LM agents and review defense countermeasures in Sect. IV-C. Finally, other security threats to LM agents are identified in Sect. IV-D, including false and harmful content generation, DoS attacks, and agent hijacking attacks.
Figure 15: The taxonomy of security threats to LM agents.
IV-A Hallucination
@@
The hallucination of LM agents typically refer to erroneous, inaccurate or illogical outputs that deviate from user inputs, generated context, or real-world conditions. It poses a substantial risk to the reliability of LM agents. For example, hallucination may lead to erroneous decision-making in an automated driving system, thereby elevating the potential for severe traffic accidents. According to [33], we categorize the hallucination within the context of LM agents from the following four perspectives, as illustrated in Fig. 16.
@@
  • Input-conflicting hallucination: The input-conflicting hallucination refers to the content generated by LM agents diverges from user input. For example, when a user requests an LM agent to draft an introduction about electric vehicles, the agent provides an introduction about gas-powered vehicles instead.
@@
  • Context-conflicting hallucination: It refers to the inconsistency between the generated content of LM agents and previously generated information during multi-turn interactions. For example, a user and an agent discuss the film “Jurassic Park” in the initial stage of a dialogue, but the agent’s responses may shift the subject to “Titanic” as interaction continues, thereby resulting in a context inconsistency.
@@
  • Knowledge-conflicting hallucination: LM agents generally depend on plug-in knowledge databases to facilitate accurate and efficient responses. The occurrence of knowledge-conflicting hallucination is noted when the responses generated by agents contradict the corresponding knowledge within knowledge databases.
@@
  • Fact-conflicting hallucination: The fact-conflicting hallucination arises when LM agents generate content which is in conflict with the established world facts. For example, a history-education LM agent provides the incorrect year 1783 as the founding date of the United States, conflating the end of the American Revolutionary War with the actual date of the nation’s founding.
@@
The causes of hallucinations of LM agents can be broadly attributed into the following three sources: data and knowledge, training and inference, and multi-agent interactions.
IV-A1 Hallucination from Data and Knowledge

### v2

@@ III-F Cooperation Paradigms of Connected LM Agents
Model output fusion Directly aggregate the outputs of different models. [118]
Knowledge Fusion KG completion Encode text to better use textual & semantic infomation. [119]
KG verification Using a small LM to validate and correct the LM output. [120]
Knowledge Retrieval Static knowledge base Retrieve knowledge from static sources for accurate reasoning. [58, 59]
Dynamic knowledge base
Use real-time updated knowledge sources for accurate reasoning.
@@
c) Knowledge retrieval. LM agents not only rely on the learned knowledge during pre-training but also can dynamically access and query external knowledge bases (e.g., databases, the Internet, and KGs) to obtain the latest information to help reasoning. RAG technology combines information retrieval and generation models by first retrieving relevant contents and then generating answers based on these contents. According to the data type, it can be divided into two categories.
  • RAG based on static knowledge sources, such as Wikipedia and documents. Lewis et al. [58] demonstrate how RAG generates accurate and contextually relevant responses by retrieving relevant knowledge chunks from static sources. For retrieval enhancement, Trivedi et al. [59] propose a new multi-step QA method named IRCoT to interleave retrieval with steps in CoT, which first uses CoT to guide retrieval and then uses retrieval results to improve CoT.
@@
  • RAG based on dynamic knowledge sources, such as news APIs. It contains three lines: exploring new knowledge, retrieving past knowledge, and self-guided retrieval. (i) For new knowledge exploration, Dai et al. [80] use RAG technology for improved safety and reliability of autonomous driving systems by utilizing real-time updated data sources, including in-vehicle sensor data, traffic information, and other driving-related dynamic data. It demonstrates RAG’s potential in handling complex environments and responding to unexpected situations. (ii) For past knowledge retrieval, Kuroki et al. [78] develop a novel vector database named coordination skill database to efficiently retrieve past memories in multi-agent scenarios to adapt to new cooperation missions. (iii) For self-guided retrieval, Wang et al. [77] propose a method called self-knowledge-guided retrieval (SKGR), which harnesses both internal and external knowledge for enhanced retrieval by allowing LLMs to adaptively call external resources when handling new problems.
III-G Summary and Lessons Learned
LM agent networks bridge the human, physical, and cyber worlds, with an architecture that integrates cloud, edge, and end devices. Powered by advanced technologies including foundation models, knowledge-related technologies, interaction technologies, multi-agent collaboration, and digital twin technologies, LM agent networks enable seamless communication, sophisticated reasoning, and scalable task execution. By leveraging hierarchical frameworks, semantic-aware communication, information-centric routing, and spatiotemporal awareness, LM agent networks optimize resource allocation, enhance real-time decision-making, and improve both intra- and inter-agent communication.
  • Intra-agent communication allows different components of a single LM agent, such as planning, action, memory, and security modules, to function cohesively. Frameworks such as the pub/sub mechanism facilitate efficient data exchange based on relevance. Inter-agent communication enables multiple LM agents to collaborate, share resources, and coordinate actions for collective intelligence. Protocols such as FIPA ACL and KQML define the vocabulary, message structure, and interaction strategies essential for effective communication between agents. Communication protocols should align with network environments (e.g., wireless or wired), latency requirements, robustness, and security needs. Adapting these protocols to various systems and agent platforms presents a significant challenge, highlighting the need for flexible, interoperable solutions.
  • The cloud-edge-end architecture allows for scalable and efficient task allocation, addressing the varying computational capabilities and real-time requirements of different devices.
  • Horizontal, vertical, and hybrid collaborations provide complementary solutions to complex problems, combining distributed and sequential task processing to optimize overall system performance. This includes cooperative LM fine-tuning, where agents collaboratively adjust LMs using domain-specific data; cooperative LM inference, where inference tasks are distributed across agents to speed up processing; and collaborative edge LM caching, where LM (or LM parts) are stored and shared across edge devices to reduce latency and improve access efficiency.
  • The synchronization of implicit and explicit knowledge is crucial for enhancing the adaptability and reliability of LM agents, enabling them to respond effectively in dynamic, evolving environments. Implicit knowledge, often derived from experience and context, allows agents to quickly adapt to unforeseen situations, while explicit knowledge, i.e., structured, defined information, ensures agents can make informed decisions based on clear rules or facts.
Figure 20: The taxonomy of security threats to LM agent networks.
IV Security Threats & Countermeasures to Large Model Agent Networks
In previous sections, we have explored the core concepts of LM agents and the importance of their networks for efficient collaboration and data/knowledge sharing. However, the three-tiered decentralized nature of LM agent networks, while offering scalability and flexibility, presents significant challenges related to secure and reliable collaboration throughout the life cycle of LM agent services. In this section, we present a comprehensive review of security threats related to LM agent networks and examine the state-of-the-art countermeasures to defend against them. Fig. 20 illustrates the taxonomy of security threats to LM agent networks.
Figure 21: Illustration of authentication threats in LM agent networks: (a) user-agent authentication, (b) agent-agent authentication threats, and (c) agent-TPSP (third-party service provider) authentication.
IV-A Authentication Threats in LM Agent Networks
Efficient and reliable authentication mechanisms is a prerequisite for safeguarding the functionality and integrity of LM agent systems. Based on the interacting entities, authentication threats to LM agents can be categorized into three types: user-agent authentication threats, agent-agent authentication threats, and agent-TPSP (third-party service provider) authentication threats, as illustrated in Fig. 21.
IV-A1 User-Agent Authentication Threats
LM agents generally store a wealth of private data for users and extensive sensitive corporate information for enterprises, making them prime targets for adversaries. Attackers can exploit identity forgery, credential theft, authentication bypassing, or privilege abuse to gain unauthorized access, leading to data breaches, system misuse, or service disruptions. Unlike traditional authentication scenarios, adversaries can leverage the advanced generative capabilities of LM agents to enhance their ability to bypass authentication. For instance, as illustrated in Fig. 21(a), an attacker could deploy an AIGC agent to forge a digital identity, imitating a legitimate user’s voice, appearance, and behavior using publicly available media from social platforms to evade multi-factor authentication (MFA). If successful, the attacker could exploit the vast personal data within LM agents to construct a digital replica of the victim or access critical corporate assets, resulting in greater losses than traditional authentication threats. Furthermore, as privatized LM agents become increasingly prevalent, interaction contexts are likely to contain even more personal and sensitive user information. This amplifies the impact of authentication threats, potentially leading to severe privacy violations and significant damage.
IV-A2 Agent-Agent Authentication Threats
As LM agents increasingly collaborate to tackle complex tasks through horizontal, vertical, or hybrid coordination, authentication between LM agents becomes critical to LM agent networks. A primary threat in such scenarios is agent spoofing attacks, where adversaries compromise or impersonate legitimate agents within the collaboration workflow. For instance, in horizontal or vertical collaboration scenarios, attackers may compromise LM agents, converting them into malicious entities, or impersonate legitimate agents to disrupt collaborative LM inference processes, thereby undermining the accuracy and reliability of the collaboration system and jeopardizing the task’s outcome, as illustrated in Fig. 21(b). Additionally, collusive LM agents can sabotage the consensus-building process required for task execution, introducing misinformation or bias that disrupts the overall functionality and trustworthiness of the collaborative system.
IV-A3 Agent-TPSP Authentication Threats
LM agents often rely on external tools, web APIs, and third-party service providers (TPSPs) to execute diverse tasks, making secure and reliable authentication between agents and TPSPs critical. This process involves mutual verification: (i) TPSPs should confirm the agent’s legitimacy and authorization, typically via a valid API key or token; (ii) agents should authenticate TPSPs to ensure they interact with trusted and legitimate service endpoints. Authentication threats in this context primarily target the communication channel between LM agents and TPSPs, exposing them to interception-based attacks such as man-in-the-middle (MITM) and session hijacking. For instance, an adversary could intercept agent-TPSP communications, altering responses or injecting erroneous data, leading to unexpected outputs for users, as illustrated in Fig. 21(c). Additionally, improper API key or token management can exacerbate risks and result in unauthorized service misuse.
IV-A4 Countermeasures to Authentication Threats in LM Agent Networks
While existing technologies such as end-to-end encryption, certificates, MFA, and key management are effective in mitigating authentication threats in LM agent services, their effectiveness diminishes in the face of the growing complexity of LM agent collaboration and the evolving sophistication of attacks driven by LMs. Moreover, the rapid development of LM agents has created a dual-edged dynamic, where both attacks and defenses are increasingly automated, escalating the risks and complexities of maintaining security. As a result, it is crucial to develop adaptive and collaborative resilient authentication mechanisms to ensure the security and reliability of LM agent systems in complex, dynamic environments.
  • Risk-Based Adaptive Authentication: This approach leverages the contextual reasoning capabilities of LM agents to dynamically assess the risk of user interactions. By analyzing user behavior, intent, and contextual anomalies, the authentication risk of the entity being authenticated is evaluated in real-time. Based on the risk level, the authenticator (such as the user, LM agent, or TPSP) adapts the authentication requirements, which may include triggering additional verification steps or involving third-party trusted institutions. This adaptive and flexible mechanism enhances the responsiveness of LM agents to varying levels of risk.
  • Multi-Agent Cooperative Authentication: This approach is designed for scenarios requiring long-term collaboration between LM agents. For instance, when multiple LM agents are working together to complete a task, each agent independently verifies the outputs or requests of the target agent for potential anomalies. The output is integrated into the collaborative workflow only after a sufficient number of agents confirm its reliability, based on a predefined threshold. This cooperative verification process reduces the risks of single-point failures, strengthens the robustness of authentication in dynamic agent networks, and mitigates the likelihood of adversarial manipulation.
Figure 22: Illustration of DoS atttacks in LM agent networks: (a) sponge examples and (b) intelligent botnet.
IV-B Denial-of-Service (DoS) Attacks in LM Agent Networks
The fine-tuning and inference processes of LM agents require significant computational resources. DoS attacks, however, can drastically increase resource consumption, overwhelming LM agents and disrupting service availability. Based on the attack method, DoS attacks in LM agent networks can be classified into two types: data-oriented DoS attacks (e.g., sponge examples) and flooding-oriented DoS attacks (e.g., intelligent botnets), as illustrated in Fig. 22.
IV-B1 Sponge Examples
Sponge examples are adversarially crafted inputs that can significantly increase the energy consumption and computational latency of neural networks. Shumailov et al. [121] demonstrate how sponge examples can be used to execute DoS attacks on AI services, causing latency and energy consumption to rise by a factor of 30. LM agents are particularly vulnerable to such attacks, as adversaries can leverage LMs to generate large volumes of optimized sponge examples at a relatively low cost, thereby amplifying the impact and making the attack more difficult to mitigate.
IV-B2 Intelligent Botnet
A botnet refers to a network of compromised Internet-connected devices controlled by adversaries, which can be exploited for various malicious activities, including distributed DoS attacks. Intelligent botnets take advantage of advanced AI technologies, particularly in the context of LMs, to enhance their attack capabilities. Unlike traditional botnets, intelligent botnets utilize the comprehension abilities of LMs to analyze vulnerabilities within target systems, autonomously refine attack strategies, and optimize their methods for greater impact. For instance, an intelligent botnet may analyze a multi-agent collaborative system to identify the most critical agent in the workflow, then launch a flooding request attack on this agent, ultimately causing a collapse of the entire system.
IV-B3 Countermeasures to DoS Attacks in LM Agent Networks
Countermeasures to DoS attacks can be categorized into two main types: passive defense and active defense. Passive defense focuses on detecting and filtering malicious inputs, while active defense is exemplified by the use of honeypot technology.
  • Malicious inputs detection and filtering technologies can effectively identify abnormal inputs, such as sponge examples, by comparing their energy consumption with that of natural inputs [121]. Specifically, prior to deploying LM agents, an analysis of natural inputs can be conducted to measure the time or energy consumption during inference. The defender can then establish a cutoff threshold to limit the maximum energy consumption for each inference, thus mitigating the impact of malicious inputs (e.g., sponge examples).
  • Honeypot technologies can be employed in a multi-agent collaborative system by deploying a specific LM agent as an intelligent honeypot to lure adversaries and divert DoS attacks away from critical agents. The attacks directed at the honeypot agent would not affect the overall system’s availability. Additionally, through security traceability and behavioral analysis, defenders can study adversarial attack patterns and enhance the system’s defense mechanisms against future threats.
IV-C Hallucination in LM Agent Networks
@@
Decision-making reliability is crucial for practically deploying LM agent services. The hallucination phenomenon in LM agents refers to the generation of erroneous or illogical outputs that deviate from user inputs, generated context, or real-world conditions, posing a significant risk to service reliability. In multi-agent scenarios, hallucinations can be exacerbated through agent interactions, amplifying and propagating errors across the system. According to [34], we categorize the hallucination within the context of LM agents from the following four perspectives, as illustrated in Fig. 23.
@@
  • Input-conflicting hallucination: The input-conflicting hallucination refers to the content generated by LM agents diverges from user input. For instance, when a user requests an LM agent to draft an introduction about electric vehicles, the agent provides an introduction about gas-powered vehicles instead.
@@
  • Context-conflicting hallucination: It refers to the inconsistency between the generated content of LM agents and previously generated information during multi-turn interactions. For instance, a user and an agent discuss the film “Jurassic Park” in the initial stage of a dialogue, but the agent’s responses may shift the subject to “Titanic” as interaction continues, thereby resulting in a context inconsistency.
@@
  • Knowledge-conflicting hallucination: LM agents generally depend on plug-in knowledge bases to facilitate accurate and efficient responses. The occurrence of knowledge-conflicting hallucination is noted when the responses generated by agents contradict the corresponding knowledge within knowledge bases.
@@
  • Fact-conflicting hallucination: The fact-conflicting hallucination arises when LM agents generate content which is in conflict with the grounded world facts. For instance, a history-education LM agent provides the incorrect year 1783 as the founding date of the United States, conflating the end of the American Revolutionary War with the actual date of the nation’s founding.
@@
The causes of hallucinations of LM agents can be broadly attributed into the following three sources: data & knowledge, training & inference, and multi-agent interactions.
IV-C1 Hallucination from Data and Knowledge